### v1

# A Point-Based Algorithm for Distributional Reinforcement Learning in Partially Observable Domains

Abstract
In many real-world planning tasks, agents must tackle uncertainty about the environment’s state and variability in the outcomes of any chosen policy. We address both forms of uncertainty as a first step toward safer algorithms in partially observable settings. Specifically, we extend Distributional Reinforcement Learning (DistRL)—which models the entire return distribution for fully observable domains—to Partially Observable Markov Decision Processes (POMDPs), allowing an agent to learn the distribution of returns for each conditional plan. Concretely, we introduce new distributional Bellman operators for partial observability and prove their convergence under the supremum p-Wasserstein metric. We also propose a finite representation of these return distributions via $\psi$-vectors, generalizing the classical $\alpha$-vectors in POMDP solvers. Building on this, we develop Distributional Point-Based Value Iteration (DPBVI), which integrates $\psi$-vectors into a standard point-based backup procedure—bridging DistRL and POMDP planning. By tracking return distributions, DPBVI naturally enables risk-sensitive control in domains where rare, high-impact events must be carefully managed. We provide source code to foster further research in robust decision-making under partial observability.
1 Introduction
Reinforcement Learning (RL) traditionally maximizes expected returns, treating each future reward distribution as a scalar. However, growing evidence shows that modeling the full return distribution, rather than just its mean, can yield robust policies, better exploration, and richer theoretical insights (Bellemare et al., 2023). This distributional perspective, referred to as Distributional Reinforcement Learning (DistRL), captures variability and higher-order statistics of returns, providing agents with more profound information about environment stochasticity.
While DistRL has been well-studied in fully observable Markov Decision Processes (MDPs), many real-world systems are partially observable (Chiam et al., 2024; Komorowski et al., 2018; Caballero-Martin et al., 2024), forcing the agent to infer the latent state from noisy or incomplete observations. Such problems are naturally framed as Partially Observable Markov Decision Processes (POMDPs), but the literature on distributional methods in these settings remains sparse.
Beyond maximizing expected returns, many real-world settings require risk-sensitive planning, where limiting the probability of adverse or catastrophic outcomes is crucial. In partially observable domains, the stakes are even higher because the agent’s uncertainty about the state can exacerbate the impact of rare but high-cost events. For instance, a slight misjudgment of the patient’s condition or the vehicle’s surroundings in healthcare or autonomous navigation can lead to undesirable or dangerous consequences. Incorporating risk considerations into POMDP planning explicitly accounts for such worst-case scenarios, making policies more conservative when high losses are possible and balancing exploration against the need to avoid significant harm. Establishing a foundational framework for distributional methods in POMDPs, as we do here, is thus a necessary first step toward risk-sensitive control in partially observable domains.
In this work, we present a formal extension of DistRL to POMDPs, bridging the gap between distributional theory and partial observability. Concretely, we make three major contributions:
  • Formal Extension of DistRL to POMDPs. We define the partially observable distributional evaluation operator $\widetilde{\mathcal{T}}_{PO}$ and its optimality operator $\widetilde{\mathcal{T}}^{*}_{PO}$, showing that both are $\gamma$-contraction under the supremum p-Wasserstein metric ($1\leq p<\infty$). This result generalizes classical DistRL convergence results to the partially observable setting.
  • Finite Representation via $\psi$-Vectors. We introduce $\psi$-vectors, a distributional analog to the well-known $\alpha$-vectors in POMDP theory. In the risk-neutral regime, we show that a finite set of $\psi$-vectors suffices to represent the optimal distributional value function while preserving the piecewise linear and convex (PWLC) property in the Wasserstein space.
  • Distributional Point-Based Value Iteration (DPBVI). We adapt Point-Based Value Iteration (PBVI) to the distributional setting, yielding DPBVI, which uses $\psi$-vectors instead of $\alpha$-vectors. Under risk-neutral objectives, DPBVI converges to the same policy as PBVI yet provides a foundation for risk-sensitive extensions by maintaining full return distributions.
Alongside these theoretical results, we release our source code [^1] Omitted for double-blind review. to facilitate further exploration of risk-aware approaches in partially observable domains. Thus, our work unifies partial observability and DistRL, enabling distributionally robust decision-making where state uncertainty and return variability are central considerations.
2 Setting
We consider a finite, discrete-time Partially Observable Markov Decision Process (POMDP) $\mathcal{P}$. POMDPs are defined as a tuple: $\langle\mathcal{S,A,O},T,\Omega,b_{0},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ is a set of discrete states, $\mathcal{A}$ is a set of discrete actions, $\mathcal{O}$ is the discrete space of noisy and/or incomplete state information, $T(s,a,s^{\prime})=P(s_{t+1}=s^{\prime}\mid s_{t}=s,a_{t}=a)$ is the state transition model, $\Omega(o^{\prime},s^{\prime},a)=P(o_{t+1}=o^{\prime}\mid s_{t+1}=s^{\prime},a_{t}=a)$ is the sensor model, $b_{0}\in\Delta$ is the initial belief, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, and $10\leq\gamma<$ 1 is the discount factor.
In this setting, the agent is unable to directly observe the state of the environment and must rely on a developed belief state $b\in\Delta$, a probability distribution across $\mathcal{S}$. The belief state serves as sufficient information for the agent to behave optimally. The agent’s belief is developed using the sequence of observations $b_{t}=P(s_{t}\mid b_{0},a_{0},o_{1},\dots,o_{t-1},a_{t-1},o_{t})$ by
$b_{t}(s_{t})$ $=\tau(b_{t-1},a,o)$ (1)
$=\dfrac{\Omega(o_{t},s_{t},a_{t-1})\sum_{s_{t-1}\in\mathcal{S}}T(s_{t-1},a_{t-1},s_{t})b_{t-1}(s_{t-1})}{\sum_{s_{t}\in\mathcal{S}}\Omega(o_{t},s_{t},a_{t-1})\sum_{s_{t-1}\in\mathcal{S}}T(s_{t-1},a_{t-1},s_{t})b_{t-1}(s_{t-1})}$
In POMDPs, the objective is to learn an optimal, stationary policy $\pi^{*}(a\mid b)$ which maximizes the expected return $\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}\mathcal{R}(s_{t},a_{t})\right]$ for time horizon $T$. Policies in this setting may be viewed as conditional plans and are typically learned by learning the optimal value function
$V^{*}(b)=\max_{a\in\mathcal{A}}\left[\sum_{s\in\mathcal{S}}\mathcal{R}(s,a)b(s)+\gamma\sum_{o^{\prime}\in\mathcal{O}}\sum_{s^{\prime}\in\mathcal{S}}\Omega(o^{\prime},s^{\prime},a)\sum_{s\in\mathcal{S}}T(s,a,s^{\prime})b(s)V^{*}(b^{\prime})\right]$ (2)
The $n$-th horizon value function is comprised of a set of $\alpha$-vectors $V_{n}=\{\alpha_{0},\alpha_{1},\dots,\alpha_{m}\}$ and is piecewise linear and convex (PWLC) (Sondik, 1978). Each $\alpha$-vector is a $|\mathcal{S}|$-dimensional hyperplane and defines the value function for some bounded region of $\Delta$ (i.e., $\max_{\alpha\in V}\alpha\cdot b$). At each planning step, the next value function $V_{n}$ may be computed from the previous value function $V_{n-1}$ via the Bellman optimality operator $\mathcal{T}_{PO}$
$V_{n}(b)$ $=\mathcal{T}_{PO}V_{n-1}$ (3)
$=\max_{a\in\mathcal{A}}\left[\sum_{s\in\mathcal{S}}\mathcal{R}(s,a)b(s)+\gamma\sum_{o^{\prime}\in\mathcal{O}}\max_{\alpha\in V_{n-1}}\sum_{s^{\prime}\in\mathcal{S}}\Omega(o^{\prime},s^{\prime},a)\sum_{s\in\mathcal{S}}T(s,a,s^{\prime})b(s)\alpha(s^{\prime})\right]$
2.1 Point-Based Value Iteration
PBVI computes $V_{t}=\mathcal{T}_{PO}V_{t-1}$ in three steps. First, it creates projections for each action and observation
$\Gamma^{a,*}\leftarrow\alpha^{a,*}(s)=\mathcal{R}(s,a)$ (4)
$\Gamma^{a,o^{\prime}}\leftarrow\alpha_{i}^{a,o^{\prime}}(s)=\gamma\sum_{s^{\prime}\in\mathcal{S}}T(s,a,s^{\prime})\Omega(o^{\prime},s^{\prime},a)\alpha_{i}^{t-1}(s^{\prime}),\forall\alpha_{i}^{t-1}\in V_{t-1}$
Next, the value of action $a$ at belief $b$ is computed by
$\Gamma_{b}^{a}=\Gamma^{a,*}+\sum_{o^{\prime}\in\mathcal{O}}\operatorname*{arg\,max}_{\alpha\in\Gamma^{a,o^{\prime}}}(\alpha\cdot b)$ (5)
Lastly, the best action at each belief is used to construct $V_{t}$
$V_{t}\leftarrow\operatorname*{arg\,max}_{\Gamma_{b}^{a},\forall a\in\mathcal{A}}(\Gamma_{b}^{a}\cdot b),\forall b\in\mathcal{B}$ (6)
$V_{t}$ is comprised of at most $|\mathcal{B}|$ $\alpha$-vectors, thus requiring $|\mathcal{S}||\mathcal{A}||V_{t-1}||\mathcal{O}||\mathcal{B}|$ operations per backup.
2.2 Distributional Reinforcement Learning
Traditionally, the value function models the return expectation. DistRL, instead, aims to learn the compound distribution of the returns $Z^{\pi}(s)$ sourced to randomness in 1. reward $R$ 2. transition $P^{\pi}$, and 3. distribution of the next-state value $Z(S^{\prime})$. Let $\widetilde{\mathcal{T}}$ be the distributional Bellman operator such that
[…]
The distributional Bellman operator has been shown to be a $\gamma$-contraction under the supremum $p$-Wasserstein metric space $\forall p\in[1,\infty]$ (Bellemare et al., 2023). The distributional Bellman optimality operator $\widetilde{\mathcal{T}}^{*}$, which selects actions that maximize the expected return, similarly converges to the fixed point under the supremum $p$-Wasserstein metric space $\forall p\in[1,\infty]$ (Bellemare et al., 2023) if there is a unique optimal policy and a mean-preserving distribution representation $\mathcal{F}$ is used (e.g., a categorical representation). In the case of the existence of multiple optimal policies, however, convergence isn’t guaranteed because differing distributions can be similar in expectation.
3 Distributional Reinforcement Learning Under Partial Observability
In the presence of partial observability we must account for additional sources of randomness; namely, randomness in (a) the unobservable state $\mathcal{S}$, (b) observations $O$ and (c) sensing $\Omega^{\pi}$. Therefore, the return distribution is learned with respect to belief instead of state (i.e., $Z(b)$).
3.1 The Partially Observable Distributional Bellman Operators
Let $\widetilde{\mathcal{T}}_{PO}$ be the partially observable distributional Bellman operator such that
[…]
Proposition 1.
For any POMDP $\mathcal{M}$ with rewards bounded by $[R_{\text{min}},R_{\text{max}}]$ the partially observable distributional Bellman operator is a contraction on $\mathfrak{P}(\mathbb{R})^{\Delta}$ in the supremum p-Wasserstein distance $\bar{w}_{p},\forall p\in[1,\infty),\forall\gamma\in[0,1)$. That is,
$\bar{w}_{p}(\widetilde{\mathcal{T}}^{\pi}_{PO}\eta,\widetilde{\mathcal{T}}^{\pi}_{PO}\eta^{\prime})\leq\gamma\bar{w}_{p}(\eta,\eta^{\prime}),$
$\forall\eta,\eta^{\prime}\in\mathfrak{P}(\mathbb{R})^{\mathcal{S}}$.
Proof.
To prove Proposition 1, we lift $\mathcal{M}$ to an equivalent, fully observable belief MDP $\widetilde{\mathcal{M}}$, where belief transitions are defined by $\tau(b,a,o^{\prime})$ with probability $P(o^{\prime}\mid b,a)=\sum_{s,s^{\prime}\in\mathcal{S}}b(s)\,T(s,a,s^{\prime})\,\Omega(o^{\prime}\mid s^{\prime},a)$. Let $p\in[1,\infty),\gamma\in[0,1)$ be fixed.
$w_{p}^{p}\left((\widetilde{\mathcal{T}}^{\pi}_{PO}Z_{0})(b),(\widetilde{\mathcal{T}}^{\pi}_{PO}Z_{1})(b)\right)$ $\overset{(a)}{\leq}\mathbb{E}_{\pi}\left[|(R+\gamma Z_{0}(B^{\prime}))-(R+\gamma Z_{1}(B^{\prime}))|^{p}\mid B=b\right]$
=$\overset{(b)}{=}\gamma^{p}\mathbb{E}\left[|Z_{0}(B^{\prime})-Z_{1}(B^{\prime})|^{p}\mid B=b\right]$
$\overset{(c)}{\leq}\gamma^{p}\int_{b^{\prime}\in\Delta}P^{\pi}(B^{\prime}=b^{\prime}\mid B=b)\mathbb{E}\left[|Z_{0}(b^{\prime})-Z_{1}(b^{\prime})|^{p}\right]$
$\overset{(d)}{\leq}\sup_{b^{\prime}\in\Delta}\mathbb{E}\left[|Z_{0}(b^{\prime})-Z_{1}(b^{\prime})|^{p}\right]$
=$\overset{(e)}{=}\gamma^{p}\bar{w}^{p}_{p}(Z_{0},Z_{1})$
(a) follows from the Wasserstein distance being defined as a minimum over couplings, (b) results from algebraically adjusting the expectation, (c) follows from the fact that the belief transitions and the return random variables are treated independently in the coupling, and the belief space being continuous necessitates integration, (d) results from the maximum of nonnegative quantities being at least as great as their weighted average, and (e) follows since $(Z_{0},Z_{1})$ is defined as an optimal coupling. ∎
Proposition 1 follows from the standard $\gamma$-contraction properties of $\widetilde{\mathcal{T}}^{\pi}$. Because $\widetilde{\mathcal{T}}^{\pi}_{PO}$ is a contraction mapping, by the Banach Fixed-Point Theorem, $\exists!\eta^{*}\in\mathfrak{P}$ s.t. $\eta^{*}$ is a fixed point.
Proposition 2.
Let $\widetilde{\mathcal{T}}^{*}_{PO}$ be the partially observable distributional Bellman optimality operator for a POMDP $\mathcal{M}$ with bounded rewards and let $p\in[1,\infty]$. Suppose there is a unique optimal policy $\pi^{*}$ and, for each state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, the reward distribution is $P_{\mathfrak{R}}(\cdot\mid s,a)$. For any initial return-distribution function $\eta_{0}\in\mathfrak{P}(\mathbb{R})$, the sequence of iterates
$\eta_{k+1}\;=\;\widetilde{\mathcal{T}}^{*}_{PO}\,\eta_{k}$
converges in the supremum $p$-Wasserstein distance $\overline{w}_{p}$ to $\eta^{\pi^{*}}$, the return distribution associated with the unique optimal policy.
Proof.
We again view the POMDP $\mathcal{M}$ through its equivalent fully observable belief MDP $\widetilde{\mathcal{M}}$, where each belief $b\in\Delta$ in $\mathcal{M}$ becomes a state of $\widetilde{\mathcal{M}}$. Noting this setup is directly analogous to the greedy selection rule $\mathcal{G}$ in Bellemare et al. (2023, Theorem 7.9), we can directly apply this theorem and conclude that the iterates converge to the unique, optimal return distribution $\eta^{\pi^{*}}$. ∎
Corollary 1.
Consider a mean-preserving projection $\Pi_{\mathcal{F}}$ for some representation $\mathcal{F}$. Suppose $\Pi_{\mathcal{F}}$ is a nonexpansion in the supremum $p$-Wasserstein distance, and let the finite domain $\mathfrak{P}_{\mathcal{R}}(\mathbb{R})$ be closed under $\Pi_{\mathcal{F}}$. If there is a unique optimal policy $\pi^{*}$ then, under the conditions of Proposition 2, the sequence
$\eta_{k+1}=\Pi_{\mathcal{F}}\widetilde{\mathcal{T}}^{*}_{PO}\eta_{k}$
converges to the fixed point $\eta^{\pi^{*}}$.
Proposition 2 states if there is a unique optimal policy, repeatedly applying $\widetilde{\mathcal{T}}^{*}_{PO}$ will converge to the optimal return distribution. Corollary 1 establishes that if we approximate the return distribution after each greedy update, the resulting approximate iteration still converges to the same fixed point, provided there is a unique optimal policy.
3.2 $\psi$-vectors
In classical POMDP value iteration, each conditional plan—a sequence of actions contingent on future observations—is captured by an $\alpha$-vector, which stores said plan’s expected return for each state. Thus, a belief-state’s value is found by selecting the $\alpha$-vector with the highest dot product $\alpha\cdot b$. While this approach gives the expected returns of each plan, it does not capture the uncertainty around possible outcomes.
To address this, we adopt a distributional viewpoint: rather than storing a scalar expectation, we store a distribution over returns for each plan. We call these $\psi$-vectors. Formally, a $\psi$-vector is
$\Psi=\{\psi^{s}\mid s\in\mathcal{S}\},$ (9)
where $\psi^{s}$ is a distribution over returns conditioned on being in state s. Given a set of candidate $\psi$-vectors $\Gamma$, the distributional value function at a belief $b$ is then obtained by taking the maximum distributional inner product:
$Z(b)\;=\;\max_{a\in\mathcal{A}}\max_{\Psi\in\Gamma_{a}}\;\big{\langle}\Psi,b\big{\rangle},$ (10)
where $\Gamma_{a}\subset\Gamma$ is the subset of $\psi$-vectors associated with action a. Here, $\langle\Psi,b\rangle$ denotes a mixture distribution of the ${\psi^{s}}$, weighted by the belief $b$. In the risk-neutral setting, we typically evaluate this mixture by taking its expectation, mirroring the $\alpha\cdot b$ calculation from classical POMDPs—but retaining the entire distribution allows us to capture richer information about potential outcomes.
Proposition 3.
Under the assumptions of Proposition 2 and Corollary 1, the optimal distributional value function $Z^{*}(b)$ can be represented as a finite set of $\psi$-vectors $\Psi$, $\forall a\in\mathcal{A}$, such that for any belief state $b\in\Delta$,
$Z^{*}(b)=\max_{a\in\mathcal{A}}\max_{\Psi\in\Psi^{*}_{a}}\langle\Psi,b\rangle,$
and these $\psi$-vectors maintain piecewise linearity and convexity within the Wasserstein metric space.
Proof.
By corollary 1, there exists a unique optimal distributional value function $Z^{*}$. Let $\Gamma^{*}=\{\Psi^{*}_{a}\mid\forall a\in\mathcal{A}\}$. Then for any belief $b$,
$Z^{*}(b)$ $=\max_{a\in\mathcal{A}}\max_{\Psi\in\Psi^{*}_{a}}\langle\Psi,b\rangle$
=$\overset{(a)}{=}\max_{\Psi\in\Gamma^{*}}\langle\Psi,b\rangle$
=$\overset{(b)}{=}\max_{\alpha\in\Gamma^{*}}\alpha\cdot b$
=$\overset{(c)}{=}V^{*}(b)$
(a) follows from the definition of $\Gamma^{*}$ as the union of $\Psi_{a}$ for all $a$, (b) because of risk-neutrality, $\max_{\Psi\in\Gamma^{*}}\langle\Psi,b\rangle$ is determined by the maximum expectation. That is, $\operatorname*{arg\,max}_{\Psi}\langle\Psi,b\rangle=\operatorname*{arg\,max}_{\Psi}\mathbb{E}[\langle\Psi,b\rangle]$, where $\alpha=\mathbb{E}[\Psi]$. (c) follows since $\alpha\in\Gamma^{*}$ precisely matches the expected returns that define the classical optimal value function $V^{*}$ in a POMDP. Finally, by Sondik’s result (Sondik, 1978), $V^{*}$ is PWLC and has finite representation. Therefore, $Z^{*}$ is PWLC and may be represented by a finite set of $\psi$-vectors. ∎
By Proposition 3, $Z^{*}$ inherits the PWLC structure of classical POMDPs, but within the Wasserstein metric space. Crucially, this representation remains finite in the risk-neutral case, ensuring computational tractability—just as a finite set of $\alpha$-vectors suffices to represent the classical optimal value function $V^{*}$.
4 Distributional Point-Based Value Iteration
Building upon the foundations of PBVI and DistRL in the partially observable setting, we now introduce Distributional Point-Based Value (DPBVI). Conceptually, DPBVI is PBVI with $\widetilde{\mathcal{T}}^{*}_{PO}$ in place of $\mathcal{T}^{*}_{PO}$. As in standard PBVI, we retain point-based belief expansion (Pineau et al., 2003), but our backup step updates a finite set of $\psi$-vectors rather than $\alpha$-vectors.
Recall that in PBVI, we maintain a finite set of belief points $\mathcal{B}\subset\Delta$, where $\Delta$ is the probability simplex over states $\mathcal{S}$. Each backup step creates new $\alpha$-vectors to approximate the PWLC value function, then selects one $\alpha$-vector per belief. DPBVI extends this idea to the distributional setting by replacing $\alpha$-vectors with $\psi$-vectors, where each $\psi$-vector stores one distribution per state for a given action.
By Proposition 3, a finite set of these $\psi$-vectors suffices to represent the optimal distributional value function $Z^{*}$ in the risk-neutral setting, just as $\alpha$-vectors suffice in classical POMDPs. We represent return distributions using the categorical distribution parameterization, which is mean-preserving when its support spans all possible returns (Bellemare et al., 2023). This ensures that the expected returns align with the classical POMDP value function, preserving the key convergence guarantees from Corollary 1 while operating in the space of return distributions.
4.1 The DBPVI Backup
Given the previous set of $\psi$-vectors $\Gamma^{t-1}$, each DPBVI backup follows four steps: 1. action-observation projection, 2. categorical distribution projection, 3. belief-specific maximization, and 4. candidate $\psi$-vector selection.
We create candidate $\psi$-vectors by combining the immediate rewards $\widetilde{R}_{a}$ with the next-step distributions of returns, conditioned on actions $a$ and observations $o^{\prime}$. Formally,
$\Gamma_{a}^{o^{\prime},t}\;\leftarrow\;\Bigl{\{}\widetilde{R}_{a}+\gamma\,T(\cdot,a,s^{\prime})\,\Omega(o^{\prime},s^{\prime},a)\,\psi^{s^{\prime}}\,\Big{|}\;\Psi\in\Gamma^{t-1},\,s^{\prime}\in\mathcal{S}\Bigr{\}}.$ (11)
This step yields $|\mathcal{A}|\times|\mathcal{O}|\times|\Gamma^{t-1}|$ projected distributions, denoted by $\Gamma^{,t}$. Because each distribution reflects one action-observation pair, it may be subnormalized until we later sum over observations.
The categorical distribution representation is not closed under $\widetilde{\mathcal{T}}^{*}_{PO}$ due to the $\gamma$-scaling and immediate reward addition operations on prior $\psi$-vectors. This means the subnormalized distributions in $\Gamma^{,t}$ may have heterogeneous supports. To ensure each $\psi$-vector has the same support, we apply the categorical projection operator $\Pi_{c}$
$\Gamma^{,t}\leftarrow\{\Pi_{c}(\Gamma_{a}^{o^{\prime},t})\mid\Gamma_{a}^{o^{\prime},t}\in\Gamma^{,t}\}.$
For each belief $b\in\mathcal{B}$, we choose the combination of $\psi$-vectors across observations that yields the highest expected distribution (under the risk-neutral lens). Symbolically, for a single action $a$,
$\Gamma_{a}^{b,t}\;=\;\sum_{o^{\prime}\in\mathcal{O}}\operatorname*{arg\,max}_{\Psi\in\Gamma_{a}^{o^{\prime},t}}\bigl{\langle}\Psi,b\bigr{\rangle}.$ (12)
This step mirrors the cross-sum and maximize operation in PBVI, except each summand is a distributional mixture rather than a scalar $\alpha-$vector. Summing those best subnormalized distributions across all observations recovers a full next-step return distribution.
Finally, for each belief $b$, we pick the best action’s distribution by
$\Psi_{b}^{t}\;\leftarrow\;\operatorname*{arg\,max}_{\Gamma_{a}^{b,t},\,a\in\mathcal{A}}\bigl{\langle}\Gamma_{a}^{b,t},b\bigr{\rangle}.$ (13)
Collecting $\{\psi_{b}^{t}\}_{b\in\mathcal{B}}$ yields our updated set $\Gamma^{t}$. As in PBVI, we keep at most $|\mathcal{B}|$ vectors in $\Gamma^{t}$, thus maintaining a point-based approximation.
Because we manage distributions instead of scalars, each backup is slower by a factor of $|M|$, where $|M|$ is the number of bins in each categorical distribution. Concretely, one DPBVI backup runs in $\mathcal{O}(|\mathcal{S}|\;|\mathcal{A}|\;|\Gamma^{t-1}|\;|\mathcal{O}|\;|\mathcal{B}|\;|M|)$, compared to $\mathcal{O}(|\mathcal{S}|\;|\mathcal{A}|\;|V_{t-1}|\;|\mathcal{O}|\;|\mathcal{B}|)$ for PBVI. Nonetheless, DPBVI preserves the core PBVI structure—belief set, backup iteration, and finite policy representation—while enabling distributional modeling of returns under patial observability.
5 Experimental Results
Our primary goal in these experiments is to verify that DPBVI converges to the same solution as PBVI under risk-neutral objectives, rather than to improve upon the performance or efficiency of PBVI. Indeed, DPBVI is slower due to the additional cost of handling distributions, but it offers a framework that can be extended to risk-sensitive criteria in future work.
5.1 Environments
We evaluate on two small POMDP environments, each with its own dynamics models:
MiniGrid DoorKey (5x5)
This environment, adapted from the MiniGruid suite (Chevalier-Boisvert et al., 2023, MiniGrid-DoorKey-5x5-v0), places an agent in a 5x5 grid containing a locked door and a key. The agent’s objective is to reach a goal state behind a locked door requiring a key. We assume both perfect sensor and transition models, a discount factor $\gamma=0.9$, a reward of $1.0$ upon reaching the goal state, and a known start state. Because the start state is known and the dynamics models are perfect, we set $\mathcal{B}$ to include complete confidence in being in one of the ten states along the optimal path (i.e., $|\mathcal{B}|=10$).
Two-State Noisy-Sensor
This is a simple domain with two states, $\{s_{0},s_{1}\}$, and a noisy sensor adapted from the environment in Russell & Norvig (2020, Section 17.5). The two actions are Go and Stay. Go changes states with probability $0.9$ and Stay stays in the same state with probability $0.9$. The agent received a reward of $1.0$ each timestep it is in stay $s_{1}$ and 00 otherwise. With probability $0.6$ the sensor reports the correct state. We set the discount factor $\gamma=0.99$ and $\mathcal{B}$ to include twenty evenly spaced beliefs between complete belief of being in either state (i.e., $|\mathcal{B}|=20$).
5.2 Setup
We disable belief-point expansion in both environments and fix a finite set $\mathcal{B}$ of belief states as described in Section 5.1. PBVI is implemented as in (Pineau et al., 2003), while DPBVI follows the backup procedure from Section 4. Return distributions in DPBVI use 51 categorical atoms, with each $\psi$-vector initialized to place all probability mass at zero. We fix the convergence criterion to $\epsilon=1\mathrm{e}{-3}$.
For the MiniGrid DoorKey environment, we set the distribution support to $[0,5]$, reflecting the small range of cumulative rewards in a 5×5 grid. In the Two-State Noisy-Sensor environment, we adopt a larger support $[0,100]$ to accommodate higher returns. Upon convergence, we compare the expected values of DPBVI’s distributions against the scalar value function from PBVI, using identical discount factors, reward functions, and stopping criteria in both algorithms.
5.3 Results and Discussion
Table 1 compares the average runtimes and the number of iterations required for DPBVI and PBVI to converge on the MiniGrid DoorKey and Two-State environments. In both domains, DPBVI recovers the same optimal risk-neutral solution as PBVI, but the added cost of handling full return distributions significantly increases its runtime.
Table 1: Results of DPBVI and PBVI: average runtime and iteration counts. Both methods converge to the same value function.
Environment Algorithm Avg. Runtime (s) # Iterations
MiniGrid DoorKey PBVI 0.014 11
DPBVI 54.805 11
Two-State PBVI 0.083 789
DPBVI 0.397 789
In MiniGrid DoorKey, PBVI completes in $\approx 0.014$ seconds on average, whereas DPBVI requires $\approx 54.8$ seconds, though both converge in 11 iterations. For the Two-State problem, DPBVI is also slower ($0.3966$s vs. $0.0825$s), despite both methods converging in 789 iterations. These patterns illustrate how distributional operations can become computationally expensive, yet DPBVI preserves the optimal solution under risk-neutral objectives and paves the way for future risk-sensitive extensions.
6 Conclusion
In this paper, we presented the first comprehensive study of Distributional Reinforcement Learning in partially observable domains. Our main theoretical result showed that the partially observable distributional Bellman operator $\widetilde{\mathcal{T}}_{PO}^{\pi}$ and its optimality counterpart $\widetilde{\mathcal{T}}^{*}_{PO}$ are $\gamma$-contractions in the supremum p-Wasserstein metric ($1\leq p<\infty$). These results extend the well-known contraction properties of distributional Bellman operators in fully observable settings to POMDPs. As part of our analysis, we introduced $\psi$-vectors as the distributional analogs of $\alpha$-vectors, proving that under risk-neutral control, the optimal distributional value function in a POMDP can be represented by a finite collection of these $\psi$-vectors while preserving PWLC in the Wasserstein metric space. This representation underlines how a distributional perspective naturally generalizes classical POMDP theory.
Building on these foundations, we described DPBVI, a risk-neutral algorithm that applies the distributional Bellman optimality operator to a finite set of belief points. This extension parallels the classical PBVI procedure yet replaces scalar $\alpha$-vectors with distributional $\psi$-vectors, thereby capturing richer information about uncertainty in returns. We showed that DPBVI recovers the same solution as PBVI when taking expectations, confirming that the distributional approach does not compromise solution quality in the risk-neutral regime. While our primary focus was on risk-neutrality, this work provides a foundation for risk-sensitive approaches to partial observability, which may require new techniques and theoretical insights to handle more general utility functions and objectives. We hope these developments foster further exploration of DistRL in practical domains where risk and uncertainty are fundamental to decision-making.